# Beyond Autoregression: Discrete Diffusion for Complex Reasoning and Planning

### v1

@@ Appendix B Algorithms for Training and Inference
Algorithm 2 Sampling from MDM
@@ C.1 Task Details
C.2 MDM Implementation Details
We conduct all the experiments on NVIDIA V100-32G GPUs, and we use 8 GPUs for training and sampling. We mainly consider comparing diffusion and AR models trained from scratch with different model sizes, with arguments for each size listed in Table 5. We use the GPT-2 architecture for both MDM and AR. We set the learning rate to 1e-3 for the tiny model and 3e-4 for others, and we set the batch size to 1024 across all the models and tasks. We train MDM for 1200 epochs on the minimal planning task, 300 epochs on Sudoku, and 600 epochs on other datasets. By default, we set the diffusion sampling steps to $T=20$ for tasks with average output tokens larger than 20, otherwise $T=10$. We use a decoding temperature $\tau=0.5$ for all tasks. For all the experiments, we have verified the statistical significance by running them multiple times.
@@ C.3 Baseline Implementation Details
We train the AR model until convergence, and the number of training epochs is set to 200 for the minimal planning task, 300 for SAT, and 40 for others. We keep other parameters, e.g., batch size and learning rate, the same as training MDM.
@@
For all the diffusion baselines, we use the same transformer architecture as GPT-2 to control the variables. We set the training parameters the same as MDM, e.g., number of training epochs to 600, learning rate to 3e-4, and batch size to 1024. During inference, we set decoding timesteps to 20 for all diffusion models as we didn’t observe a clear performance improvement when scaling to 1024.
@@ D.1 Token Consumption on Game of 24
MDM 76.0 11 26
D.2 Case Study
@@
Numbers Goal Groundtruth AR Prediction MDM Prediction

### v2

@@ Appendix B Algorithms for Training and Inference
Algorithm 2 Sampling from MGDM
@@ C.1 Task Details
C.2 MGDM Implementation Details
We conduct all the experiments on NVIDIA V100-32G GPUs, and we use 8 GPUs for training and sampling. We mainly consider comparing diffusion and AR models trained from scratch with different model sizes, with arguments for each size listed in Table 5. We use the GPT-2 architecture for both MGDM and AR. We set the learning rate to 1e-3 for the tiny model and 3e-4 for others, and we set the batch size to 1024 across all the models and tasks. We train MGDM for 1200 epochs on the minimal planning task, 300 epochs on Sudoku, and 600 epochs on other datasets. By default, we set the diffusion sampling steps to $T=20$ for tasks with average output tokens larger than 20, otherwise $T=10$. We use a decoding temperature $\tau=0.5$ for all tasks. For all the experiments, we have verified the statistical significance by running them multiple times.
@@ C.3 Baseline Implementation Details
We train the AR model until convergence, and the number of training epochs is set to 200 for the minimal planning task, 300 for SAT, and 40 for others. We keep other parameters, e.g., batch size and learning rate, the same as training MGDM.
@@
For all the diffusion baselines, we use the same transformer architecture as GPT-2 to control the variables. We set the training parameters the same as MGDM, e.g., number of training epochs to 600, learning rate to 3e-4, and batch size to 1024. During inference, we set decoding timesteps to 20 for all diffusion models as we didn’t observe a clear performance improvement when scaling to 1024.
@@ D.1 Token Consumption on Game of 24
MGDM 76.0 11 26
D.2 AR with token reweighting
We show the accuracy of AR with the same token reweighting mechanism in Equation 8 on the minimal planning task in Table 7. We find that applying token reweighting to AR models still cannot solve subgoals with PD larger than 1 (i.e., with accuracy around 50%), similar to the original AR.
Table 7: Results of AR with token reweighting.
Planning Distance AR AR with token reweighting
0 100 100
1 100 100
2 51.1 52.1
3 46.9 51.5
4 52.0 50.3
5 49.9 51.9
D.3 Scaling both Data and Model Size
As an extension of Table 1, we show the accuracy of AR and MGDM when both data and model size are increased in Table 8. We find scaling both data and model size is effective for both AR and MGDM.
Table 8: Results of scaling both model and data size.
AR MGDM
85M model, 500k data 45.8 91.5
303M model, 500k data 41.3 88.3
303M model, 1M data 53.3 95.6
D.4 Case Study
@@
Numbers Goal Groundtruth AR Prediction MGDM Prediction